### v1

# Multi-Agent Path Finding under Limited Communication Range Constraint via Dynamic Leading

Abstract
This paper proposes a novel framework to handle a multi-agent path finding problem under a limited communication range constraint, where all agents must have a connected communication channel to the rest of the team. Many existing approaches to multi-agent path finding (e.g., leader-follower platooning) overcome computational challenges of planning in this domain by planning one agent at a time in a fixed order. However, fixed leader-follower approaches can become stuck during planning, limiting their practical utility in dense-clutter environments. To overcome this limitation, we develop dynamic leading multi-agent path finding, which allows for dynamic reselection of the leading agent during path planning whenever progress cannot be made. The experiments show the efficiency of our framework, which can handle up to 25 agents with more than 90% success-rate across five environment types where baselines routinely fail.
Introduction
We want a team of agents navigate through an obstacle-rich environment to goals while maintaining constant team communication: a spanning tree created from range-limited communication between pairs of agents. This problem is relevant to scenarios like supply delivery during disasters or monitoring hostile environments, where the risk of losing agents is significant. To mitigate this risk, agents must ensure that the team is in constant communication throughout their movement. Maintaining the spanning tree while the agents head in different directions with varied lengths of actions makes pathfinding challenging in continuous time and space, even for holonomic agents. The challenge is compounded by agents starting at random positions, needing to pass through narrow passages and non-convex spaces without collisions, and then reach random goals.
The problem can theoretically be solved using a centralized approach [1, 2, 3, 4, 5], in which planning selects between team actions, each simultaneously specifying an action for all agents at once. However, this approach quickly suffers from the curse of dimensionality, as the difficulty of planning increases exponentially with the number of agents, making planning intractable for even relatively small problems. The platooning leader-follower approach [6, 7, 8, 9] mitigates this challenge by planning each agent in sequence, where one agent is selected as the leader and others act as followers, who plan so as to maintain communication to the agent that planned before them. However, establishing a fixed planning order for leader agent can result in issues during planning. These issues include the follower agents finding no action to follow the leader within communication range or becoming stuck once the leader reaches its goal (Fig. 1.a), or the team needing to spread out for the followers to reach their goals while maintaining communication range (Fig. 1.b).
Figure 1: Fixed leader-follower approaches fail in scenarios a) & b). Our framework, MA-DL, can handle both (c). If $a_{1}$’s leading causes the team to get stuck (a), dynamic leading allows another agent take over as the new leader. When the leader and follower move to different directions (b), followers are allowed to pursue another agent to its goal.
To address these challenges, this work develops a framework for solving multi-agent pathfinding problem with limited communication range (MALCR), ensuring a connected communication channel in the team at any time. We propose a novel technique: dynamic leading which allows any agent to become the new leader during multi-agent tree expansion, allowing planning to further expand the multi-agent tree if the current leader cannot make progress. Once the leader agent has planned, follower agents plan, ensuring that the resulting path is always in communication to at least one agent.
We introduce an algorithm for planning in multi-agent systems with communication distance constraints: MA-DL. The experiments show that our framework results in fast and effective planning, handling up to 25 agents with more than 90% success-rate across five environment types where baselines, centralized planning with composite states and platooning leader-follower approach, routinely fail.
Figure 2: An illustration of one step expanding the MATree with 3 agents. $a_{1}$, $a_{2}$, and $a_{3}$, start at $s_{1}$, $s_{2}$, and $s_{3}$, aiming for their goals $g_{1}$, $g_{2}$, and $g_{3}$ while avoiding the obstacles (black) and other agents. From node $v_{1}$ in the multi-agent tree $\mathcal{T}_{ma}$ (step 1), the planner implements single plannings with random planning order (step 2–5), then expand $\mathcal{T}_{ma}$ based on the paths found (step 6). In step 3, $a_{2}$ follows $a_{3}$ to maintain communication links. At $t_{1}$(F), when $a_{3}$ stops, $a_{2}$ switches to follow $a_{1}$. Our key contribution, dynamic leading, occurs at step 4 when $a_{3}$ follows $a_{2}$ to $t_{2}$ (L), then becomes the new leader and plan to $t_{3}$. In the next iteration, other agents follows $a_{3}$ before expanding the MATree in step 6.
Related Work
There exist multiple approaches that seek to solve multi-agent path finding problems, in the absence of a communication constraint, that could in theory be adapted to solve the MALCR problem. To effectively scale to more agents, approaches in this space are typically designed around reducing the exponential growth of the planning tree with planning time horizon. Wagner and Choset [1] proposed M algorithm, which plans individually for each agent, yet builds composite states for agents and expands the local planning tree when their paths are in collision. Though M works well if the agents are far apart, when agents are close to each other, collisions occurs frequently, causing the planning joint-state increases rapidly until fully incooperating of all agents and limiting its efficacy for solving our limited-communication problem, which requires that agents be near to one another. Solovey, Salzman, and Halperin [2] introduced the dRRT planner, which uses a implicit composite roadmap—a product of single-agent roadmaps. Their algorithm grows a multi-agent tree in this composite roadmap using RRT. However, using such a roadmap requires abstracting low-level details of each robot’s motion over time, making it difficult to apply their approach for planning with communication constraints. PIBT [5] and BMAA* [3] expand a single action for each agent in each planning iteration. These planners can handle well deadlocks among the agents. However, there are two issues for these planners to solve MALCR problem. The first one is their short-term action selection strategy, which can result in situations where followers are unable to find actions that keep them within the leader’s communication range. The second one is their fixed planning order causing followers fail to plan when their initial positions are out of the leader’s communication range (Fig. 1b). Okumuar et al. [4] proposed the OTIMAPP planner which employs Deadlock-based Search (DBS) to resolve conflict among paths determined through prioritized planning. In addition to the limitation of a fixed planning order (as mentioned above), OTIMAPP requires a new technique to modify the computed paths while adhering to deadlock and communication constraints. In general, it is very challenging to adapt these approaches to solve the MALCR problem in which each member of the team must be in constant communication with one another in continuous time.
Specific to the MALCR problem, platooning has emerged as a state-of-the-art solution for multi-agent path finding in which the motion of a leader-agent is planned first and follower-agents planned one at a time in sequence to follow the agent that proceeded them, and so is designed to maintain communication constraints between pairs of agents. Most existing work in platooning performs full motion planning only for the leader, and uses a low-level controller to regulate followers to maintain a distance from the preceding agents [7, 6]. Qian et al. [8], Zhao et al. [10], and Gao et al. [9] introduced model predictive controller that generate trajectories for a virtual center, which followers need to track.
However, recent work in the space of conflict-based search (CBS) [11] has proven that planning with a fixed vehicle order in general, including platooning, is incomplete and frequently becomes stuck when the leader moves in a direction different to the follower’s goals, as we show in Fig. 1, or when the team must shuffle their planning order during travel to reach their goals without violating communication constraints. Overcoming this limitation in general requires a planner that allows for dynamic leading, in which the leader can change during multi-agent tree expansion when the team fails to make progress towards the goal.
Problem Definition
We formally define the Multi-Agent Path Finding with Limited Communication Range (MALCR) problem as follows. There are $n$ agents $\{a_{1},\ldots,a_{n}\}$ and a known world $\mathcal{W}$ with obstacles $\mathcal{O}=\{\mathcal{O}_{1},\ldots,\mathcal{O}_{m}\}$. The agents start at the initial positions $s^{{\mathrm{init}}}={\langle{s^{{\mathrm{init}}}_{1},\ldots,s^{{\mathrm{init}}}_{n}}\rangle}$ and head to the goals $g=\{g_{1},\ldots,g_{n}\}$, where $s_{i},g_{i}\in\mathcal{W}$ and $s_{i},g_{i}\notin\mathcal{O}$. The initial positions and goals are chosen so as to satisfy the team communication constraint, formally defined below.
The world $\mathcal{W}$ is divided into sub-divisions $\Delta$, which are obstacle-free regions. A graph $\mathcal{G}$ is created where the vertices are the centroids of $\Delta$ and the edges represent connections of neighboring sub-divisions. An agent moves to a neighboring sub-division with constant velocity ${{\mathrm{v}}}_{c}$ by taking the action: Move$(v,u)$, where $v,u$ are neighbor vertices in $\mathcal{G}$. As reaching its goal, the agent takes the action of Stop = Move$(v,v)$. Agents moving over $\mathcal{G}$ are guaranteed to be collision-free with respect to the static obstacles. Time and agents’ positions are continuous due to varied lengths of actions.
We define two levels of communication constraint for the agents. The first level is between two agents, referred to as agent communication constraint (ACOMM), which requires the distance between their positions to be less than or equal to the communication range $r_{c}$. The second level applies to the entire agent team, referred to as team communication constraint (TCOMM), which requires the team to form a spanning tree where the edges represent the connections between pairs of agents satisfying the ACOMM constraint. An action Move($v,u$) satisfies the ACOMM constraint if during the movement Move($v,u$) the agent has at least one neighboring agent within a distance of $r_{c}$.
A collision occurs between two agents, $a_{i}$ and $a_{j}$, when their distance is less than a threshold $d_{c}$ at timestep $t$. A path is a sequence of waypoints to transition an agent from a position $p_{s}$ to position $p_{g}$ with constant velocity $\emph{v}_{c}$. We say a path is reach-goal if it can lead the agent to the goal, is collision-free, and the movement between two sequential waypoints satisfies ACOMM constraint; When an agent stops at its goal, we still consider collision and ACOMM constraints. A path is valid if it is reach-goal and at the goal $p_{g}$ from $t_{g}$, the agent’s action Stop still satisfies ACOMM and is collision-free.
The objective is to compute valid paths $\{\zeta_{1},\ldots,\zeta_{n}\}$, one for each agent, so that $\zeta_{i}$ starts at $s^{{\mathrm{init}}}_{i}$, reaches $g_{i}$, and the agents satisfy TCOMM constraints from the start time to the time the last agent reaches its goal. We develop a MAPF framework that seeks to reduce the overall planning time and the travel distances.
Algorithm 1 MA-DL module
  INPUT: $n$: number of agents, $g=\{g_{1},...,g_{n}\}$: goals, $\mathcal{W}$: ${{\mathrm{world}}}$, $s^{{\mathrm{init}}}=\{s_{1}^{{\mathrm{init}}},...,s_{n}^{{\mathrm{init}}}\}$: starts, $t_{{\mathrm{ds}}}$: max runtime
  OUTPUT: paths $\zeta=\{\zeta_{1},...,\zeta_{{\mathrm{n}}}\}$
1:  $\mathcal{T}_{{\mathrm{ma}}}\!\leftarrow{\scriptstyle{{\textsc{InitMATree}}}}(s^{{\mathrm{init}}},t=0)$; $\Delta\!\leftarrow\!{\scriptstyle{{\textsc{SubDivision}}}}(\mathcal{W})$; $\mathcal{G}\leftarrow{\scriptstyle{{\textsc{CreateGraph}}}}(\Delta,g)$
2:  $\mathcal{H}\leftarrow\{{\scriptstyle{{\textsc{CalHeursitic}}}}(\mathcal{G},g_{1}),\ldots,{\scriptstyle{{\textsc{CalHeursitic}}}}(\mathcal{G},g_{n})$}
3:  while ${\scriptstyle{{\textsc{time}}}}()<t_{{\mathrm{ds}}}$ do
4:     $v\leftarrow{\scriptstyle{{\textsc{SelectNode}}}}(\mathcal{T}_{{\mathrm{ma}}})$; $p\leftarrow{\scriptstyle{{\textsc{InitOrder}}}}(v)$; $\zeta^{s}\leftarrow{\scriptstyle{{\textsc{InitPaths}}}}(v)$
5:     if ($v.{{\mathrm{visited}}}$) then ${\scriptstyle{{\textsc{Shuffle}}}}(p)$
6:     for $1\leq k\leq m$ do
7:        ${{\mathrm{allRG}}}\leftarrow{{\texttt{true}}}$; if $k>m_{t}$ then ${\scriptstyle{{\textsc{Shuffle}}}}(p)$
8:        for $1\leq i\leq n$ do
9:           if ${\scriptstyle{{\textsc{FoundGoal}}}}(\zeta^{s}_{p_{i}})$ then continue
10:           $\zeta^{t}\!\leftarrow\!{\scriptstyle{{\textsc{SAPF}}}}(p_{i},g_{p_{i}},\zeta^{s},\mathcal{G},\mathcal{H}_{i})$; $\zeta^{s}_{p_{i}}.{\scriptstyle{{\textsc{Insert}}}}(\zeta^{t})$
11:           if ${\scriptstyle{{\textsc{FoundGoal}}}}(\zeta^{s}_{p_{i}})$ then ${\scriptstyle{{\textsc{ModifyIfOverlap}}}}(p_{i},\zeta^{s},{{\mathrm{allRG}}})$
12:           else ${{\mathrm{allRG}}}={{\texttt{false}}}$
13:        if ${{\mathrm{allRG}}}$ then break
14:     ${\scriptstyle{{\textsc{ExpandMATree}}}}(\mathcal{T}_{{\mathrm{ma}}},\zeta^{s},v)$; $v.{{\mathrm{visited}}}\leftarrow{{\texttt{true}}}$
15:     if ${{\mathrm{allRG}}}$ then break;
16:  ${{\mathrm{vid}}}={\scriptstyle{{\textsc{ClosestNodeToGoal}}}}(\mathcal{T}_{{\mathrm{ma}}})$
17:  return ${\scriptstyle{{\textsc{GetPaths}}}}(\mathcal{T}_{{\mathrm{ma}}}({{\mathrm{vid}}}))$
Methods
Algorithm 2 SAPF module
  INPUT: $\rho$: agent’s ID, $g_{\rho}$: goal, $\mathcal{G}$: init. graph, $\zeta^{s}$: previous planned paths, $t_{{\mathrm{sa}}}$: max runtime, $\mathcal{H}_{\rho}$: shortest path heuristic
  OUTPUT: a path $\zeta_{\rho}$
1:   $\mathcal{G}^{s}={\scriptstyle{{\textsc{AddNodes}}}}(\mathcal{G},g_{\rho},\zeta^{s}_{\rho}.{{\mathrm{end}}})$
2:  $v_{g}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(g_{\rho},\mathcal{G}^{s})$; $v_{s}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(\zeta^{s}_{\rho}.{{\mathrm{end}}},\mathcal{G}^{s})$
3:  ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(v_{s})$; ${{\mathrm{closedSet}}}\leftarrow\{\}$; $v_{{\mathrm{best}}}\leftarrow v_{s}$
4:  while ${{\mathrm{openList}}}$ not empty and ${\scriptstyle{{\textsc{time}}}}()<{{\mathrm{t}}}_{{\mathrm{sa}}}$ do
5:     $v\leftarrow{{\mathrm{openList}}}.{\scriptstyle{{\textsc{Pop}}}}()$; ${{\mathrm{closedSet}}}.{\scriptstyle{{\textsc{Add}}}}(v)$
6:     if $v=v_{g}$ then return ${\scriptstyle{{\textsc{getpath}}}}(v_{s},v,\mathcal{G}^{s})$
7:     for $u$ in ${\scriptstyle{{\textsc{GetNeighbors}}}}(v)$ do
8:        if $u\textbf{ is not in }{{\mathrm{closedSet}}}$ then
9:           $u.{{\mathrm{parent}}}\leftarrow v$; $d_{uv}\leftarrow{\scriptstyle{{\textsc{Distance}}}}(v,u)$; $u.g\leftarrow v.g$ + $d_{uv}$
10:           $u.t\leftarrow v.t$ + $d_{uv}/\text{v}_{c}$; $u.h\leftarrow{\scriptstyle{{\textsc{GetHeuristic}}}}(u,\mathcal{H}_{\rho})$; $u.f=u.g+u.h$
[…]
12:              ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(u)$
13:              if $u=v_{g}$ then return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},u,\mathcal{G}^{s})$
14:        else if $v.g+d_{uv}<u.g$ then
15:           $u_{1}\leftarrow{\scriptstyle{{\textsc{Copy}}}}(u)$; $u_{1}.t\leftarrow v.t+d_{uv}/\text{v}_{c}$
16:           if IsNodeValid($\rho,\zeta^{s},u_{1},v,r_{c}$then
17:              $u.{{\mathrm{parent}}}\leftarrow v$; $u.t\leftarrow v.t+d_{uv}/\text{v}_{c}$; $u.g\leftarrow v.g+d_{uv}$; $u.f=u.g+u.h$
18:        if $v_{{\mathrm{best}}}.f>u.f$ then $v_{{\mathrm{best}}}\leftarrow u$
19:  return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},v_{{\mathrm{best}}},\mathcal{G}^{s})$
Our framework, named as MA-DL, consists of two modules: Multi-agent Path Finding with Dynamic Leading (MA-DL) and Single-Agent Path Finding (SAPF).
High Level Overview: The main loop of the MA-DL module (Alg. 1) expands a multi-agent (MA) planning tree $\mathcal{T}_{ma}$ whose growth is illustrated in Fig. 2. The MA tree is used to represent motion expansion of all agents. Its role is to keep track of planning progress for all agents, and the node with lowest value of cost and heuristic will be selected to expand in the next iteration. To ensure all nodes in $\mathcal{T}_{ma}$ have a chance of being selected, their costs are penalized after each selection. The MA tree expansion occurs in a loop (Fig. 2:2–5), in which each agent’s path is expanded in order. The single-agent planner (SAPF, Alg. 2) finds reach-goal shortest paths for the agents one by one according to the planning order $p$, such that each agent maintains communication with at least one of the agents that plans before it. Difference to prioritized planning approach [12] with fixed leading, the leader agent in our planner is dynamic: through dynamic leading technique (Alg. 3) which allows any agent take the leading role if it goes further than other agents, or shuffling the planning order whenener planning cannot further expand the multi-agent tree (Alg. 1:6). After each planning loop, the multi-agent tree is expanded with new nodes built from the individual paths (Fig. 2.6). Our dynamic leading technique (Fig. 2.4) allows the MA-DL planner to handle the challenging situations in Fig. 1.
MA-DL MODULE
This module manages the multi-agent tree (MATree) $\mathcal{T}_{ma}$, selects nodes to expand, dynamically selects planning orders, and triggers single-agent plannings.
The module gets the number of agents $n$, the goals $g$, initial start $s^{\text{init}}$, and the world $\mathcal{W}$ as inputs and returns the valid paths or, if valid paths cannot be found, those closest to the goal. The algorithm (Alg. 1) starts by initializing a multi-agent tree $\mathcal{T}_{ma}$ with the root consisting of all initial positions $s^{\text{init}}$ (Alg. 1:1). The world $\mathcal{W}$ is divided into sub-divisions $\Delta$ of a predefined area, with obstacle-overlapping sub-divisions further subdivided along their largest dimension until they clear or reach a size threshold. A graph $\mathcal{G}$ is created from the centroids of the obstacle-free sub-divisions, while the edges represent the connections between neighboring sub-divisions (sharing boundaries or corners). We compute the shortest paths $\mathcal{H}$ from all sub-divisions to the agent goals $g$ as heuristics by the function CalHeuristic() (Alg. 1:2). In the main loop (Alg. 1:3–15), a node $v$ with smallest cost (defined at Alg. 4.14) is selected from $\mathcal{T}_{ma}$. The cost of $v$ is then increased to encourage selecting other nodes. A heuristic function INITORDER() returns an initial planning order $p$. A list of $n$ paths $\zeta^{s}$ is initialized from $v$ (Alg. 1:4). If the node $v$ is visited, the planning order p is shuffled to get a different planning order (Alg. 1:5). The planning loop (Alg. 1:6–13) attempts to find valid paths for the agents. The loop starts by setting the variable ${{\mathrm{allRG}}}$ to true (Alg. 1:7). The planning order p is shuffled after $m_{t}$ iterations of failing to reach the goal (Alg. 1:7). The for loop (Alg. 1:8–12) plans for each agent in order by calling the module SAPF and then the returned paths is inserted into $\zeta^{s}_{p_{i}}$ (Alg. 1:10). If $\zeta^{s}_{p_{i}}$ is reach-goal, it is then checked for validity (valid) by the function ${\scriptstyle{{\textsc{MODIFYIFOVERLAP}}}}()$. The path is invalid if the Stop action at the goal at time $t$ blocks the future movement of an already-planned agent (according to planning order $p$), a situation called collision-at-goal by Bui et al. [13]. If the situation occurs, the function modifies the earlier planned paths and set ${{\mathrm{allRG}}}$ to false (Alg. 1:11). If the path is not reach-goal, ${{\mathrm{allRG}}}$ is set to false (Alg.1:12).
After all agent planning has completed, the function EXPANDMATREE() is triggered to expand $\mathcal{T}_{ma}$ from $v$ using the agent’s paths $\zeta^{s}$ (Alg. 1:14); then the node $v$ is also marked as visited. If all agents reach the goals (${{\mathrm{allRG}}}$ is still true), we break the while loop (Alg. 1:15), then return paths for all agents from $\mathcal{T}_{ma}$ (Alg. 1:16–17).
Algorithm 3 ${\scriptstyle{{\textsc{IsNodeValid}}}}(\rho,\zeta^{s},u,v,r_{c}$) (Dynamic leading is employed)
1:  ${{\mathrm{lead}}}\leftarrow{{\texttt{true}}}$; ${{\mathrm{comm}}}\leftarrow{{\texttt{false}}}$
2:  for  $1\leq i\leq|\zeta^{s}|$  do
3:     if $u.t\leq\zeta^{s}_{i}.\text{maxtime}$ then ${{\mathrm{lead}}}\leftarrow{{\texttt{false}}}$
4:  if lead then $\text{lead}\leftarrow{\scriptstyle{{\textsc{IsCOMMAtGoal}}}}(u,\zeta^{s},\rho)$
5:  for $1\leq i\leq|\zeta^{s}|$ do
6:     if ${\scriptstyle{{\textsc{IsCollision}}}}(u,v,\zeta^{s}_{i},\rho)$ then return false
7:     if ${\scriptstyle{{\textsc{IsCOMMS}}}}(u,v,\zeta^{s}_{i},\rho,r_{c})$ then ${{\mathrm{comm}}}\leftarrow{{\texttt{true}}}$
8:  return $({{\mathrm{lead}}}\ \|\ {{\mathrm{comm}}})$
SAPF MODULE
Single-Agent Path Finding (SAPF) finds a shortest reach-goal path for an agent. This planner searches actions on $\mathcal{G}$ that satisfy the ACOMM constraint and are collision-free. With enough runtime, the planner can explore all possible paths and so is probabilistically complete. If it fails, it returns the closest-to-goal path. Agents are assumed to have the same velocity ${{\mathrm{v}}}_{c}$. The inputs are the agent’s ID $\rho$, goal position $g_{\rho}$, a graph of the world sub-divisions $\mathcal{G}$, planned paths of previous agents $\zeta^{s}$, and the shortest path to the goal heuristic $\mathcal{H}$.
To initialize, a new graph $\mathcal{G}^{s}$ is created by adding to $\mathcal{G}$: (i) the current agent position $\zeta^{s}_{\rho}.{{\mathrm{end}}}$ and (ii) the goal $g_{\rho}$, each of which connect to the centroid of the sub-division to which they belong (Alg. 2:1). The variables $v_{g}$ and $v_{s}$ are the goal and current nodes (Alg. 2:2). $v_{s}$ is then added into ${{\mathrm{openList}}}$—a priority queue data structure, which sorts its elements by their cost-to-goal. closedSet and the closest-to-goal node $v_{\text{best}}$ are also initialized (Alg. 2:3).
The search loop (Alg. 2:4–18) is: a node with lowest cost is popped out from ${{\mathrm{openList}}}$ (Alg. 2:5); then we iterate through all its neighbors (Alg. 2:7–18) to find valid nodes, calculate or update their costs, parents, and timestep, then add them into openList.
Figure 3: Out-of-communication-at-goal situation (a) and how ISCOMMATGOAL() works (b). In both situations, the planning order is ($a_{3},a_{2},a_{1}$) and the leader is changed at $t_{i+2}$.
For a node $u$ to be valid, the move action from its parent $v$ must be collision-free and satisfy the ACOMM constraint. Both of these requirements are checked by the function ${\scriptstyle{{\textsc{ISNODEVALID}}}}()$ (Alg. 2:11,16).
Algorithm 4 ${\scriptstyle{{\textsc{ExpandMATree}}}}(\mathcal{T}_{{\mathrm{ma}}},\zeta^{s},v$)
1:  $v_{p}\leftarrow v$; ${{\mathrm{timeList}}}\leftarrow\{\}$ //priority queue
2:  for $0\leq$ i $<|\zeta^{s}|$  do
3:     for $0\leq$ j $<|\zeta^{s}_{i}.{{\mathrm{times}}}|$ do ${{\mathrm{timeList}}}$.${\scriptstyle{{\textsc{Insert}}}}(\zeta^{s}_{i}.{{\mathrm{times}}}[j])$
4:  for $t$ in ${{\mathrm{timeList}}}$  do
5:     $v_{n}\leftarrow{\scriptstyle{{\textsc{NewNode}}}}()$; $v_{n}({{\mathrm{parent}}},g,t)\leftarrow(v_{p},v_{p}.g,t)$
6:     for $0\leq j<|\zeta^{s}|$ do
7:        if $t>\zeta^{s}_{j}.{{\mathrm{end}}}.t$ then
8:         if ${\scriptstyle{{\textsc{FoundGoal}}}}(\zeta^{s}_{j})$ then ${{\mathrm{pos}}}\leftarrow\zeta_{j}.{{\mathrm{end.pos}}}$
9:         else return
10:        else ${{\mathrm{pos}}}\leftarrow{\scriptstyle{{\textsc{GetPosAtTime}}}}(\zeta^{s}_{j},t)$
11:        $v_{n}.{{\mathrm{state}}}.{\scriptstyle{{\textsc{Insert}}}}({{\mathrm{pos}}})$; $d={\scriptstyle{{\textsc{Dist}}}}(v_{p}.{{\mathrm{state}}}_{j},v_{n}.{{\mathrm{state}}}_{j})$
12:        $v_{n}.g$ += $d$; $\zeta^{s}_{j}$.${{\mathrm{costogoal}}}$= $d$; $v_{n}.h$ += $\zeta^{s}_{j}$.${{\mathrm{costogoal}}}$
13:     $v_{n}.f\leftarrow v_{n}.g$ + $v_{n}.h$
14:     $\mathcal{T}_{{\mathrm{ma}}}.{\scriptstyle{{\textsc{Append}}}}(v_{n})$; $v_{p}\leftarrow v_{n}$
IsNodeValid() Function
This function (Alg. 3) checks the validity of action $\textit{Move}(v,u)$ for an agent $\rho$ starting at timestep $v.t$. For agent-collision detection, ${\scriptstyle{{\textsc{IsCollision}}}}()$ (Alg. 3:6) checks if any segmented point along the line $l_{vu}$ (connecting $v$ and $u$) collides with the planned paths in $\zeta^{s}$ (Alg. 3:5) at corresponding timesteps. If a collision occurs, the node is invalid.
The ACOMM constraint check depends on whether agent $\rho$ is a leader or follower. In our approach, an agent becomes the new leader if its timestep during expansion is ahead of all others (Alg. 3:2–3); we call this dynamic leading. Because it plans farther out in time than other agents, the leader is not subject to any communication constraints when expanding its plan. A follower agent $\rho$ must satisfy the ACOMM communication constraint: that it must maintain communication with another agent while moving. The function ${\scriptstyle{{\textsc{ISCOMMS}}}}()$ (Alg. 3:7) checks that agent $\rho$ can communicate with a neighbor during its action.
During dynamic leading, a situation called out-of-communication-at-goal can occur, which is shown in Fig. 3. To prevent this situation, we propose the function IsCommAtGoal().
IsCommAtGoal() Function
In Fig. 3, at $t_{i}$ agent $a_{3}$ is the leader, and reaches its goal in two steps ($t_{i+1}$ and $t_{i+2}$). $a_{2}$ follows $a_{3}$ and also stops at its goal by $t_{i+2}$. Agent $a_{1}$ follows $a_{3}$ until $t_{i+2}$, and then becomes the leader, planning without ACOMM constraints shown in Fig. 3.a. However, the MATree $\mathcal{T}_{ma}$ expansion halts at $t_{i+6}$ due to communication breakdown. The function ${\scriptstyle{{\textsc{ISCOMMATGOAL}}}}()$ guides $a_{1}$ to remain in communication. It is triggered when the leader is changed (Alg. 3.4). It checks if any of the new leader’s neighbors have reached their destination. If so, the leader’s action $\textit{Move}(v,u)$ must meet the ACOMM constraint with an at-goal agent. If it cannot, its leader status is revoked.
EXPAND MULTI-AGENT TREE
Function ${\scriptstyle{{\textsc{ExpandMATree}}}}()$ (Alg. 4) integrates the paths from the single agents $\zeta^{s}$ and adds the combined paths to expand the multi-agent tree $\mathcal{T}_{{\mathrm{ma}}}$. Each node on $\mathcal{T}_{ma}$ has an associated timestep and the agent positions are interpolated to make them match for each single agent path.
The for loop (Alg. 4:2–3) collects all timesteps of the waypoints on the single paths and inserts into a priority queue ${{\mathrm{timeList}}}$. The second for loop (Alg. 4:4–14) goes through all elements in ${{\mathrm{timeList}}}$ to create new nodes and add valid nodes into the tree $\mathcal{T}_{ma}$. Each new node $v_{n}$ inherits the travel cost from its parent with timestep $t$ (Alg. 4:5). We then go to each agent’s path, interpolate to recover its position at time $t$ (Alg. 4:7–12), and insert it into the state of node $v_{n}$. If $t$ is larger than the max-timestep of the path $\zeta^{s}_{j}$ (Alg. 4:7) and the path reaches the goal, the final position of $\zeta^{s}_{j}$ is returned (Alg. 4:8). If the position is not the goal, the tree’s expansion stops (Alg. 4:9).
If $t$ is smaller or equal to the last timestep $\zeta^{s}_{j}$, the agent’s position at $t$ is interpolated by function ${\scriptstyle{{\textsc{GetPosAtTime}}}}()$ (Alg. 4:10). The lines in Alg. 4:11–12 add the agent’s positions into the node state; we then update the travel cost and the heuristic. Finally, the new node $v_{n}$ is added into the MATree $\mathcal{T}_{{\mathrm{ma}}}$, and $v_{p}$ is reassigned to continue the expansion (Alg. 4:14).
Experiments and Results
Experiments are conducted on five obstacle-rich environments (Fig. 4). Our planner’s performance is measured by three metrics: (1) success-rate, (2) runtime, and (3) per-agent travel distance. Our MA-DL approach is compared with two baselines: a centralized approach with composite states and a platooning leader-follower approach. All agents have 8 actions: to move one unit in the four cardinal directions and $\sqrt{2}$ unit in their diagonal directions. We also evaluate the planner’s robustness versus of runtime, goal configuration, and difficulty of environments.
Figure 4: Success-rate, Runtime, and Travel Distances of MA-DL (ours), PLF, and COMP on 5 environment types. The comm. range is 15 m and max-runtime is 5 s. The runtime and travel distances are shown by means and standard deviations.
Experimental Setup
Baselines
We have implemented two baselines: Centralized approach with composite state (COMP) and Platooning Leader-Follower approach (PLF). COMP is selected because existing MAPF planners—which typically advance the motion of each agent in isolation of one another and then attempt to resolve conflicts as necessary—are not well suited to handle the communication constraint, as the challenge of ensuring that all agents are in constant communication would require an undue number of repair operations, causing such planners to struggle (Section 2). Meanwhile, PLF is chosen because it is a state-of-the-art approach to solve MALCR problem.
  • COMP grows a multi-agent tree via A* using informed by a heuristic of sum of shortest-to-goal paths from each agent. In joint-actions of single and diagonal moves, the diagonal ones are trimmed down to equal length with the single moves.
  • PLF also builds a multi-agent tree and using priority planning to grows the tree. At the root of tree expansion, the planning order is shuffled and so a random leader is chosen; planning order is then fixed for downstream nodes. The followers plan to follow another agent if its preceding agent reached its goal. If the current expansion fails after some iterations, the planning starts again at the root with a new random leader and corresponding planning order.
Environments and Instances
We evaluate MA-DL in five obstacle-rich environment types: Random Forest, Office, Waves, Rings, and Maze (Fig. 4). All environments are square shapes of size $114\times 114$ m. On each environment type, there are one hundred maps generated with random locations of obstacles.
For each number of agents, one instance is generated on each environment map. So, there are total of 12000 instances for the experiments. The start and goal configurations are chosen randomly on alternate sides of the maps (except the ring environment), distributed with a rectangular area. The configurations are satisfied the TCOMM constraint. Communication between two agents are established if the distance between them is within 15 m of one another.
Env. Type 1: Random Forest
(Fig. 4) Obstacles with random shapes and sizes are distributed randomly occupying 10% of the environment’s area.
Env. Type 2: Office
(Fig. 4) The environments consists multiple rooms and hallways. The room has a fixed width and varying length from 9–13 m. There are three long hallways along the building, 2–3 short hallways connecting them. The hallway width varies from 7–9 m.
Env. Type 3: Waves
(Fig. 4) The environments have wave-like obstacles which are separated at regular distances. Gaps are placed in each wave with random widths. The number of waves is set to $10$.
Env. Type 4: Rings
(Fig. 4) The environments are featured by concentric rings with six breaks of random widths within 6–8 m. The separation between the rings is set to 8 m. For each instance, the starts are randomly placed at the center, and the goals are on one of four corners of the maps.
Env. Type 5: Maze
(Fig. 4) The environments consist of mazes generated by Kruskal’s algorithm with size of $14\times 14$. To ease in generation of the starting and goal configurations, boundary walls connecting to the top-most and bottom-most rows are removed.
Measuring Performance
We evaluate our framework against COMP and PLF baselines on 100 instances for each environment type and number of agent, measuring success-rate, average runtime and average distance traveled per agent. Planning is considered successful if all agents reach their goals within 5 s of runtime. Runs that exceed that planning ceiling are assigned a travel distance of 300 m. The framework is evaluated with variations over number of agents, goal configurations, runtime, and environment difficulty levels.
Computing Resources
The experiments ran on HOPPER, a computing cluster provided by GMU’s Office of Research Computing. Each planning instance is run single threaded, yet experiment were run in parallel across 48 cores on a 2.40GHz processor. Our code was developed in C++ and compiled with g++-9.3.0.
Results
Results when varying the agent number
We tested the MA-DL planner with 2–25 agents, allocating 5 seconds of runtime, and the results are shown in Fig. 4. MA-DL performs well with up to 25 agents, achieving over 90% success-rate in all environments except Maze Env. The long narrow passages of the maze mean that if some agents stop at the beginning of these passages, it becomes very difficult for others to pass through, leading to extended runtime for expanding the MATree.
Both baseline planners struggle as the number of agents is increased. PLF manages up to 5 agents in Rings Env. and only 3-4 agents in other environments. The heuristic of shortest-to-goal path sum guides COMP well to handle up 4 agents in Random Forest, Rings, and Waves Env. However, with greater than 6 agents, the high dimensionality of the search space means that planning cannot reach the goal in the allotted time and success rate quickly declines.
In addition to average success rate, we also report the mean and standard deviation for runtime and travel distance in Fig. 4. If a planner fails, the runtime and traveled distance penalties are applied to the results. The runtime and distance results aligned with the success-rate results across all planners.
Results when varying runtime
In this experiment, we run the planners with various runtime from 1 s to 512 s. We select the Maze Env., the most challenging Env. type for the experiments. The results are shown in Fig. 5.a, showing that success-rate reaches 100% as the runtime grows.
Refer to caption Refer to caption
a) Success-rate vs Runtime b) Success-rate vs Env. Difficulty
Figure 5: Performance of MA-DL as varying the runtime (a) and environment difficulty (b) with 23 agents.
Results when varying environment difficulty
To demonstrate the robustness of our framework, we run MA-DL on variations of the Rings environment type with three difficulty levels: easy, medium, and hard. The difficulty is controlled by the number of concentric rings, distance between consecutive rings, and number of breaks/gates on the rings. The features controlling difficulty are shown in Table I. An easy and hard environment are shown in Fig. 5.b.
Diff. Level Ring Count Ring-Distance Break Count
Easy 4–5 8.0 m 6–7
Medium 5 7.0 m 5–6
Hard 6 5.5 m 4–5
TABLE I: Environment Difficulty Features
The performances on three environment types with 23 agents and 5 s runtime are shown in Fig. 5.b with success-rate decreasing from easy to hard level. With medium or hard levels, MA-DL planner needs more than 5 s runtime to find valid paths for the agents.
Results with Different Goal Configurations
The goal configuration distribution impacts the planner’s performance, especially when managing diverse agent movement. A thin, long goal distribution increases runtime as the planner must search longer for a suitable planning order. We run MA-DL with an alternate goal configuration distributions: long and thin versus rectangular (Fig. 6) to test its robustness. Following with ACOMM makes the MA-DL robust in handling diverse movement directions across most environments (similar to Random Forest Env.-Fig. 6.left). However, in Maze environments with narrow passages and only one way to go, frequent collisions and path modifications significantly increase difficulty of the planning problem, reducing planning performance somewhat.
Figure 6: Long-Thin vs Rectangle of Goal Configuration.
Completeness Analysis
Though our MA-DL planner represents an advance over leader-follower planners by overcoming a limitation of fixed-priority-order that results in their incompleteness, MA-DL is also incomplete in this domain. We present an analysis of specifically what causes this incompleteness and present a direction to overcome this limitation for future work.
Ma et al. [11] proved the following theorem for the incomplete prioritized planning:
Theorem 1: Prioritized (fixed) planning with an arbitrary priority ordering is incomplete for MAPF in general.
Based on this theorem, fixed and platooning leader-follower approaches are incomplete planners. A fix leader-follower planner never changes the planning order, while a platooning planner only does so when members leave the team. However, while our MA-DL planner overcomes this limitation, it is also an incomplete algorithm. This is a consequence of the greedy nature of the single agent planner (SAPF), which always immediately returns upon finding a valid single agent path. We illustrate a scenario in Fig. 7 that MA-DL fails to solve: though the solution requires that agent $a_{3}$ select a longer path to allow other agents to pass by, the SAPF planner returns only the shortest path and so the team becomes stuck.
Future work could explore extending our MA-DL approach to achieve completeness. For this to be possible, the SAPF single-agent planner would need to be extended to eventually generate all possible paths when expanding each agent, a modification that would require careful consideration and detailed study so that it would not dramatically slow planning performance. In this research, our focus is on developing a fast and practical planner that can efficiently handle the MALCR problem, outperforming existing approaches. We plan to develop a complete version of the MA-DL planner in the future research.
Figure 7: A scenario demonstrating the incompleteness of MA-DL. Due to the greedy nature of SAPF (find the shortest path), all agents attempt to go through $G_{1}$ without trying $G_{2}$ and cannot all reach their respective goals.
Conclusion
This paper introduces the MA-DL framework to handle the MALCR problem. The core advance of our approach is dynamic leading, which enables the dynamic reselection of the leading agent during path planning whenever progress stalls, overcoming a key limitation of state-of-the-art platooning approaches. We have tested our MA-DL planner in multiple environments with features such as obstacles-rich, narrow passages, non-convex spaces, and long hallways. The results demonstrate the robustness of MA-DL planner, capable of planning up to 25 agents within 5 seconds of runtime. In future work, we aim to develop a complete version of this planner. We also intend to design a new heuristic to improve the expansion of single agent trees. Furthermore, we plan to extend the framework to support robot teams with rich kinodynamic constraints, bringing it a step closer to the real-world applications.
ACKNOWLEDGMENT
The work by E. Plaku is supported by (while serving at) the National Science Foundation. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors, and do not necessarily reflect the views of the National Science Foundation.
This material is based upon work supported by the National Science Foundation under Grant No. 2232733. G. J. Stein acknowledges support from this grant.